# Design of efficient high-order immersed metagratings using an evolutionary algorithm

### v2

@@ 3.2 Performance of the optimized metagrating
The sawtooth grating shows a smooth, gradual spectral variation compared to the metagrating’s staggered response. This smoothness is predicted from the scalar diffraction theory. According to that, an ideal blazed grating reaches peak diffraction efficiency when the operating wavelength matches the blazing wavelength. And this efficiency follows a sinc-squared envelope that rapidly declines away from the blaze [soskind2011field]. The reason for the metagrating’s staggered spectral features is not entirely clear. One possible explanation is that the discrete binary structures of the metagrating introduce sharp transitions in the electromagnetic response due to localized resonances and abrupt geometry changes between neighbouring unit cells.
@@ 3.3 Manufacturing considerations
As mentioned earlier, the feature size ($150\,\mathrm{nm}$) and the small aspect ratio ($\sim 1.3$) make this design feasible using photolithography and have been demonstrated with even smaller feature sizes ($150\,\mathrm{nm}$) [keil2021large, vakarin2021metamaterial, khaidarov2022large, hu2020cmos]. Photolithographic fabrication is inherently susceptible to systematic errors, including focus offsets, feature size variations, rounding artifacts, etch-depth inaccuracies, and mask misalignment [wang2015fabrication]. Assessing the tolerance of the proposed design to such imperfections is therefore essential. Systematic errors are typically on the order of $\sim\pm 1$$5\,\mathrm{nm}$. In the present study, the design performance was evaluated under a systematic deviation $\Delta\mathbf{x}$ from the optimal parameters $\mathbf{x}$ by computing the grating efficiency for $\mathbf{x}\pm\Delta\mathbf{x}$. The analysis was carried out until a performance degradation of approximately $10\%$ was observed, which is taken as the effective tolerance limit.
@@ 4 Summary and outlook
For this study, a more conservative estimation was made to keep the manufacturing of the designed metagrating more achievable. The restriction on the number of divisions per period (here, 5) can be relaxed and could be made a free parameter for optimization. This would be especially beneficial due to advanced e-beam lithography techniques, where sub-$10\,\mathrm{nm}$ feature sizes are possible [gour2024wafer]. Additionally, including more geometries in the optimization process—such as pillars with differently shaped bases, hollowed-out structures, and/or irregularly spaced meta-surfaces—could further help realize more control over the phase profile [so2023revisiting], which in turn would help to achieve the perfect blazed profile across a broader band over both polarizations.
It is also important to consider the limitations of the simulations. Some manufacturing defects are not included, such as imperfections in the aluminum coating/filling of the etched metasurfaces. Furthermore, wavefront errors and ghosts were not simulated, and therefore, in practice, the efficiency of the grating may slightly deviate from the simulations [zeitner2010high]. The impact of such imperfections can be best assessed by manufacturing the metagrating. Hence, the natural next step would be to test the designed metagrating by realizing it using either DUV lithography (for prototyping) or e-beam lithography (for more accurate realization).
@@
Code and Data availability The Code used for the CMA-ES optimization and the simulations of the optimized design, along with the Data underlying the results presented in this paper, are available at [patel2025_metagratingModCMAES]

### v3

@@ 3.2 Performance of the optimized metagrating
The sawtooth grating shows a smooth, gradual spectral variation compared to the metagrating’s staggered response. This smoothness is predicted from the scalar diffraction theory. According to that, an ideal blazed grating reaches peak diffraction efficiency when the operating wavelength matches the blazing wavelength. And this efficiency follows a sinc-squared envelope that rapidly declines away from the blaze [37]. The reason for the metagrating’s staggered spectral features is not entirely clear. One possible explanation is that the discrete binary structures of the metagrating introduce sharp transitions in the electromagnetic response due to localized resonances and abrupt geometry changes between neighbouring unit cells.
@@ 3.3 Manufacturing considerations
As mentioned earlier, the feature size ($150\,\mathrm{nm}$) and the small aspect ratio ($\sim 1.3$) make this design feasible using photolithography and have been demonstrated with even smaller feature sizes ($150\,\mathrm{nm}$) [22, 38, 23, 19]. Photolithographic fabrication is inherently susceptible to systematic errors, including focus offsets, feature size variations, rounding artifacts, etch-depth inaccuracies, and mask misalignment [39]. Assessing the tolerance of the proposed design to such imperfections is therefore essential. Systematic errors are typically on the order of $\sim\pm 1$$5\,\mathrm{nm}$. In the present study, the design performance was evaluated under a systematic deviation $\Delta\mathbf{x}$ from the optimal parameters $\mathbf{x}$ by computing the grating efficiency for $\mathbf{x}\pm\Delta\mathbf{x}$. The analysis was carried out until a performance degradation of approximately $10\%$ was observed, which is taken as the effective tolerance limit.
@@ 4 Summary and outlook
For this study, a more conservative estimation was made to keep the manufacturing of the designed metagrating more achievable. The restriction on the number of divisions per period (here, 5) can be relaxed and could be made a free parameter for optimization. This would be especially beneficial due to advanced e-beam lithography techniques, where sub-$10\,\mathrm{nm}$ feature sizes are possible [14]. Additionally, including more geometries in the optimization process—such as pillars with differently shaped bases, hollowed-out structures, and/or irregularly spaced meta-surfaces—could further help realize more control over the phase profile [36], which in turn would help to achieve the perfect blazed profile across a broader band over both polarizations.
It is also important to consider the limitations of the simulations. Some manufacturing defects are not included, such as imperfections in the aluminum coating/filling of the etched metasurfaces. Furthermore, wavefront errors and ghosts were not simulated, and therefore, in practice, the efficiency of the grating may slightly deviate from the simulations [41]. The impact of such imperfections can be best assessed by manufacturing the metagrating. Hence, the natural next step would be to test the designed metagrating by realizing it using either DUV lithography (for prototyping) or e-beam lithography (for more accurate realization).
@@
Data availability Data underlying the results and the code for the simulations presented in this paper are available at [30].